# The Illusion of State in State-Space Models

### v1

@@ Abstract
State-space models (SSMs) have emerged as a potential alternative architecture for building large language models (LLMs) compared to the previously ubiquitous transformer architecture. One theoretical weakness of transformers is that they cannot express certain kinds of sequential computation and state tracking (Merrill & Sabharwal, 2023a), which SSMs are explicitly designed to address via their close architectural similarity to recurrent neural networks (RNNs). But do SSMs truly have an advantage (over transformers) in expressive power for state tracking? Surprisingly, the answer is no. Our analysis reveals that the expressive power of SSMs is limited very similarly to transformers: SSMs cannot express computation outside the complexity class $\mathsf{TC}^{0}$. In particular, this means they cannot solve simple state-tracking problems like permutation composition. It follows that SSMs are provably unable to accurately track chess moves with certain notation, evaluate code, or track entities in a long narrative. To supplement our formal analysis, we report experiments showing that Mamba-style SSMs indeed struggle with state tracking. Thus, despite its recurrent formulation, the “state” in an SSM is an illusion: SSMs have similar expressiveness limitations to non-recurrent models like transformers, which may fundamentally limit their ability to solve real-world state-tracking problems.
@@ 1 Introduction
Figure 1: We prove that SSMs, like transformers, cannot solve inherently sequential problems like permutation composition ($S_{5}$), which lies at the heart of state-tracking problems like tracking chess moves in source-target notation (see Section 3.2), evaluating Python code, or entity tracking. Thus, SSMs cannot, in general, solve these problems either.
Recent theoretical work (Merrill & Sabharwal, 2023a) has shown that models built upon the transformer architecture are incapable of expressing inherently sequential computation. These results reveal a surprising limitation of transformers: they cannot express simple kinds of state tracking problems, such as composing sequences of permutations, which even simple recurrent neural networks (RNNs) can naturally express. In a different line of work, state space model (SSM) architectures (Gu et al., 2021, 2022a; Fu et al., 2023; Gu & Dao, 2023; Wang et al., 2024) have been introduced as an alternative to transformers, with the goal of achieving similar expressive power to RNNs for handling problems that are naturally stateful and sequential (Gu et al., 2021, 2022b). But does the seemingly stateful design of SSMs truly enable them to solve sequential and state-tracking problems that transformers cannot? If so, this would be a promising property of SSMs because state tracking is at the heart of large language model (LLM) capabilities such as tracking entities in a narrative (Heim, 1983; Kim & Schuster, 2023), playing chess under certain notation [^1] The hardness of chess state tracking holds with (source, target) notation, but standard notation may make state tracking easier., or evaluating code. This would motivate further research into SSM architectures and their deployment as the next generation of LLMs.
In this work, we show that the apparent stateful design of SSMs is an illusion as far as their expressive power is concerned. In contrast to the suggestion by Gu et al. (2021, 2022b) (and, perhaps, a broader belief in the community) that SSMs have expressive power for state tracking similar to RNNs, we prove theoretically that linear and Mamba-style SSMs, like transformers, cannot express inherently sequential problems, including state-tracking problems like composing permutations that RNNs can easily express. Further, our experiments confirm our formal predictions: both transformers and these SSMs cannot learn to compose permutations with a fixed number of layers, whereas RNNs can compose permutations with just a single layer. Our results imply that arguments that SSMs have an advantage over transformers due to being “more recurrent” or capable of tracking state are misguided. In fact, the SSM architectures we consider are just as theoretically unequipped for state tracking and recurrent computation as transformers are.
@@
Figure 2: Complexity hierarchy within $\mathsf{NC}^{1}$. Transformers can only recognize languages within $\mathsf{TC}^{0}$ (Merrill & Sabharwal, 2023a), and we show the same for SSMs (Theorems 4.1 and 4.2). Thus, both architectures cannot express the “hard state tracking” captured by $\mathsf{NC}^{1}$-complete problems like $S_{5}$, which can be straightforwardly expressed by RNNs. The figure assumes the widely held conjecture $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$.
Our empirical investigation shows that, in practice, both SSMs with the Mamba architecture (Gu & Dao, 2023) and transformers do not learn to solve the permutation composition state-tracking problem with a fixed number of layers, while simple RNNs can do so with just one layer. This provides empirical support for our theoretical separation in expressive power for state tracking between SSMs and true recurrent models. We also find that both transformers and SSMs struggle compared to RNNs on state-tracking problems less complex than permutation composition where it is not known whether they can express a solution. Thus, in practice, SSMs may struggle not just on the hardest state-tracking problems like permutation composition but also on easier variants.
Finally, we propose two minimal extensions of linear SSMs that increase their expressive power for state tracking, allowing them to solve permutation composition. These extensions, however, may come with a cost: they may negatively impact parallelism as well as learning dynamics. We view it as an interesting open question whether it is possible to develop SSM-like models with greater expressivity for state tracking that also have strong parallelizability and learning dynamics, or whether these different goals are fundamentally at odds, as Merrill & Sabharwal (2023a) suggest.
@@ 2 Background
We first present the SSM architectures we will analyze (Section 2.1). Our analysis of the state tracking capabilities of SSMs borrows deeply from the circuit complexity and algebraic formal language theory literature. We thus review how circuit complexity can be used to analyze the power of neural networks (Section 2.3) and how state-tracking problems can be captured algebraically and analyzed within the circuit complexity framework (Section 3.1).
@@ 2.1 Architecture of State-Space Models
SSMs consist of state-space layers, which can be thought of as simplified RNN layers. We consider two variants of the state-space layer: the linear SSM layer (of which S4 is a special case; Gu et al., 2022a) and the extended S6 layer used by Mamba (Gu & Dao, 2023).
Definition 2.1 (Linear SSM layer; e.g. S4).
Given a sequence of embeddings or previous states $x_{1},\ldots,x_{n}\in\mathbb{R}^{m}$, the recurrent form of a linear SSM layer defines a new sequence of states $h_{1},\ldots,h_{n}\in\mathbb{R}^{d}$ using learned parameter matrices $A\in\mathbb{R}^{d\times d}$ and $B\in\mathbb{R}^{d\times m}$. For each $1\leq i\leq n$,
$h_{i}=Ah_{i-1}+Bx_{i}.$ (1)
The convolutional form of the linear SSM layer defines the same [^2] The two forms express the same function over $\mathbb{R}$ or any other distributive datatype. Over floating points (Appendix A), they are not guaranteed to be the same, but we must assume the error is negligible for them to be well-defined and usable in practice. $h_{1},\ldots,h_{n}$ but is unrolled as a summation of terms:
$h_{i}=\sum_{j=1}^{i}A^{i-j}Bx_{j}.$ (2)
The layer outputs $y_{i}=Ch_{i}+Dx_{i}$ where $C\in\mathbb{R}^{d\times m}$ and $D\in\mathbb{R}^{m\times m}$.
S4 chooses a specific parameterization of a linear layer. A full S4 model is a cascade of such layers and feedforward layers, analogous to how transformers alternate multihead-self-attention layers with feedforward layers.
The S6 layer used by Mamba (Gu & Dao, 2023) generalizes a linear SSM layer by adding a selection mechanism inspired by the dynamic gating in LSTMs (Hochreiter & Schmidhuber, 1997) and GRUs (Cho et al., 2014).
Definition 2.2 (S6 layer).
An S6 layer is parameterized by a diagonal matrix $A\in\mathbb{R}^{d\times d}$, a vector $\delta\in\mathbb{R}^{d}$, and affine projections $s_{\delta}:\mathbb{R}^{m}\to\mathbb{R}^{d},s_{B}:\mathbb{R}^{m}\to\mathbb{R}^{d},$ and $s_{C}:\mathbb{R}^{d}\to\mathbb{R}^{m}$. Let $\tau$ be softplus. The S6 layer is:
$h_{i}$ $=\exp(\delta_{i}A)h_{i-1}+B_{i}x_{i},$ (3)
$\textrm{where}\;\delta_{i}$ $=\tau(\delta+s_{\delta}(x))\;\textrm{and}$
$\;B_{i}$ $=s_{B}(x).$
This implies the convolutional form:
$h_{i}=\sum_{j=1}^{i}\left(\prod_{k=j+1}^{i}\exp(\delta_{k}A)\right)B_{j}x_{j}.$ (4)
Finally, to compute $y_{i}$, $C$ is made input dependent $C_{i}$ and computed via a projection in the same manner as $B_{i}$. The layer output is then $y_{i}=C_{i}h_{i}+x_{i}$.
In practice, a crucial detail for training SSMs is the initialization of the weight matrices. Our main results (Theorems 4.1 and 4.2) will apply for any linear SSM (including S4) as well as S6, independent of the specific values of its weights. In contrast to S4 and S6, H3 (Fu et al., 2023) is not a true SSM because the context is not represented by a single vector. Rather, its architecture resembles a transformer with SSM components. Analyzing H3 is beyond our focus, but we believe our ideas could be extended to H3 in future work.
@@ 2.2 Numeric Datatype
Our main results (Theorems 4.1 and 4.2) will go through for any datatype $\mathbb{D}$ for which the following 3 operations are efficiently parallel-computable, i.e., are in the complexity class $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (to be defined shortly in Section 2.3):
@@
    Matrix powering, i.e., computing the $n$-th power of a fixed-size $k\times k$ matrix over $\mathbb{D}$
When $\mathbb{D}$ is any finite-precision datatype, i.e., has a fixed number of bits available (e.g., 16 or 64), then these operations are easily seen to be in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. As Merrill & Sabharwal (2023b) argue, however, finite-precision datatypes severely limit the expressivity of neural architectures from a formal perspective (e.g., finite-precision transformers cannot represent uniform attention), motivating the use of parameterized datatypes that can (approximately) represent any number with a sufficiently large parameter. Interestingly, when $\mathbb{D}$ is the datatype of $n$-bit integers, all of the above operations are known to be in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Hesse, 2001; Mereghetti & Palano, 2000). Realistically, however, neural model implementations use floating point numbers with much fewer than $n$ bits.
Concretely, we imagine working with the log-precision floating point model used by Merrill & Sabharwal (2023b) to analyze transformers, which we will show satisfy all these properties. For some fixed constant $c\in\mathbb{Z}^{+}$, a $c\log n$ precision float is a tuple $\langle m,e\rangle$ where $m,e$ are signed integers together taking $c\log n$ bits. Using $|x|$ to mean the number of bits used to represent integer $x$, this float represents the value $m\cdot 2^{e-|m|+1}$.
Unlike for integers, arithmetic operations over log-precision floats are not closed. That is, the product $\phi_{1}\times\phi_{2}$ of two $p$-precision floats is a well-defined number but may not be exactly representable as a $p$-precision float. It is thus necessary to define approximate versions of these operations when formalizing log-precision floating-point arithmetic. To this end, Merrill & Sabharwal (2023a) define a natural notion of approximate iterated addition over log-precision floats and show that it is computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. Following Merrill & Sabharwal (2023a)’s definition of iterated addition for floats, we define iterated multiplication and matrix powering over any datatype $\mathbb{D}$ as the result of treating the numbers as reals, performing exact arithmetic, and casting the exact output $\phi$ back to $\mathbb{D}$, denoted $\mathsf{cast}_{\mathbb{D}}(\phi)$.
In Appendix A, we extend the arguments of Hesse (2001) and Mereghetti & Palano (2000) for integers to show that iterated product and matrix powering over log-precision floats are also computable in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (cf. Lemmas 2.3 and 2.4 in Appendix A).
Lemma 2.3 (Iterated float product).
Lemma 2.4 (Float matrix power).
Let $k,c\in\mathbb{Z}^{+}$ be fixed constants. Let $M$ be a $k\times k$ matrix over $c\log n$ precision floats. Let $z\leq n,z\in\mathbb{Z}^{+}$. Then float matrix power $M^{z}$ can be computed in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$.
Combined with the result for iterated addition from Merrill & Sabharwal (2023a), this establishes that the three needed properties are met for log-precision floats.
@@ 2.3 Limits of Transformers via Circuit Complexity
A line of recent work has used circuit complexity and logic formalisms to identify the expressiveness limitations of transformers on reasoning problems (Angluin et al., 2023; Merrill & Sabharwal, 2023a; Liu et al., 2023; Chiang et al., 2023; Merrill & Sabharwal, 2023b; Hao et al., 2022); see Strobl et al., 2023 for a survey. In particular, Merrill & Sabharwal (2023a) showed that transformers can only solve problems in the complexity class $\mathsf{TC}^{0}$, which is defined as the set of problems that can be recognized by constant-depth, polynomial-size threshold circuit families. Such circuits, in addition to having standard AND, OR, and NOT gates (of arbitrary fan-in), can also use threshold gates that output $1$ iff at least $k$ of the inputs are $1$, where $k$ is a parameter of the gate. Informally, $\mathsf{TC}^{0}$ can be thought of as the class of problems that can be solved with extremely parallelized (constant-depth) computation. [^3] We use $\mathsf{TC}^{0}$ to mean $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, meaning the circuit family is constructible by a Turing machine that runs in space logarithmic in the size of the input (cf. Merrill & Sabharwal, 2023a; Strobl et al., 2023). We believe our results could be extended from $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ to $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ using techniques similar to Merrill & Sabharwal (2023b) for composing $\mathsf{TC}^{0}$ circuits in a way that preserves $\mathsf{DLOGTIME}$ uniformity.

### v2

@@ Abstract
State-space models (SSMs) have emerged as a potential alternative to transformers. One theoretical weakness of transformers is that they cannot express certain kinds of sequential computation and state tracking (Merrill & Sabharwal, 2023a), which SSMs are explicitly designed to address via their close architectural similarity to recurrent neural networks. But do SSMs truly have an advantage (over transformers) in expressive power for state tracking? Surprisingly, the answer is no. Our analysis reveals that the expressive power of S4, Mamba, and related SSMs is limited very similarly to transformers (within $\mathsf{TC}^{0}$), meaning these SSMs cannot solve simple state-tracking problems like permutation composition and consequently are provably unable to accurately track chess moves with certain notation, evaluate code, or track entities in a long narrative. To supplement our formal analysis, we report experiments showing that S4 and Mamba indeed struggle with state tracking. Thus, despite their recurrent formulation, the “state” in common SSMs is an illusion: S4, Mamba, and related models have similar expressiveness limitations to non-recurrent models like transformers, which may fundamentally limit their ability to solve real-world state-tracking problems. Moreover, we show that only a minimal change allows SSMs to express and learn state tracking, motivating the development of new, more expressive SSM architectures.
@@ 1 Introduction
Figure 1: We prove that SSMs, like transformers, cannot solve inherently sequential problems like permutation composition ($S_{5}$), which lies at the heart of state-tracking problems like tracking chess moves in source-target notation (see Section 3.2), evaluating Python code, or entity tracking. Thus, SSMs cannot, in general, solve these problems either.    Code: http://jpetty.org/ssm-illusion
Recent theoretical work has shown that transformer architecture based models are incapable of expressing inherently sequential computation (Merrill & Sabharwal, 2023a). These results reveal a surprising limitation of transformers: they cannot express simple kinds of state tracking problems, such as composing sequences of permutations, which even simple recurrent neural networks (RNNs) can naturally express. In a different line of work, state space model (SSM) architectures (Gu et al., 2021, 2022a; Fu et al., 2023; Gu & Dao, 2023; Wang et al., 2024) have been introduced as an alternative to transformers, with the goal of achieving RNN-like expressive power for handling problems that are naturally stateful and sequential (Gu et al., 2021, 2022b). But does the seemingly stateful design of SSMs truly enable them to solve sequential and state-tracking problems that transformers cannot? If so, this would be a promising property of SSMs because state tracking is at the heart of large language model (LLM) capabilities such as tracking entities in a narrative (Heim, 1983; Kim & Schuster, 2023), playing chess under certain notation [^1] The hardness of chess state tracking holds with (source, target) notation, but standard notation may make state tracking easier., or evaluating code. This would motivate further research on SSM architectures and their deployment in the next generation of LLMs.
In this work, we show that the apparent stateful design of SSMs is an illusion as far as their expressive power is concerned. In contrast to the suggestion by Gu et al. (2021, 2022b) (and, perhaps, a broader belief in the community) that SSMs have expressive power for state tracking similar to RNNs, we prove theoretically that linear and Mamba-style SSMs, like transformers, cannot express inherently sequential problems, including state-tracking problems like composing permutations that RNNs can easily express. Further, our experiments confirm this prediction: both transformers and these SSMs cannot learn to compose permutations with a fixed number of layers, whereas RNNs can compose permutations with just a single layer. Our results imply that arguments that current SSMs have an advantage over transformers due to being “more recurrent” or capable of tracking state are misguided. In fact, the SSM architectures we consider are just as theoretically unequipped for state tracking and recurrent computation as transformers are.
@@
Empirically, we find that S4 (Gu et al., 2022a) and S6 (Gu & Dao, 2023) SSMs, as well as transformers, do not learn to solve the permutation composition state-tracking problem with a fixed number of layers, while simple RNNs can do so with just one layer. This provides empirical support for our theoretical separation in expressive power for state tracking between SSMs and true recurrent models. We also find that both transformers and SSMs struggle compared to RNNs on state-tracking problems less complex than permutation composition where it is not known whether they can express a solution. Thus, in practice, SSMs may struggle not just on the hardest state-tracking problems like permutation composition but also on easier variants.
@@
Figure 2: Complexity hierarchy within $\mathsf{NC}^{1}$. Transformers can only recognize languages within $\mathsf{TC}^{0}$ (Merrill & Sabharwal, 2023a), and we show the same for SSMs (Theorems 4.2 and 4.4). Thus, both architectures cannot express the “hard state tracking” captured by $\mathsf{NC}^{1}$-complete problems like $S_{5}$, which can be straightforwardly expressed by RNNs. The figure assumes the widely held conjecture $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$.
Finally, we consider a minimal extension of a linear SSM which makes the transition matrix input dependent, similar to Liquid S4 (Hasani et al., 2023). We show that this extension has sufficient expressive power for state tracking and permutation composition. Empirically, we show that our implementation of this extension learns to solve permutation composition with a single layer, just like an RNN, while being similarly parallelizable to other SSMs. It is an open question whether such SSM architectures with greater expressivity for state tracking are practically viable for large-scale language modeling.
@@ 2 Background
We first present the SSM architectures we will analyze (Section 2.1). Our analysis of the state tracking capabilities of SSMs borrows deeply from the circuit complexity and algebraic formal language theory literature. We thus review how circuit complexity can be used to analyze the power of neural networks (Section 2.3) and how state-tracking problems can be captured algebraically and analyzed within the circuit complexity framework (Section 3).
@@ 2.1 Architecture of State-Space Models
SSMs consist of SSM layers, which can be thought of as simplified RNN layers. We define a generalized linear SSM layer that encapsulates both S4 (Gu et al., 2022a) and the S6 layer used by Mamba (Gu & Dao, 2023) as special cases.
Definition 2.1 (Generalized linear SSM layer).
Given a sequence [^2] In practice, this sequence is often a vector $\mathbf{x}_{1},\ldots\mathbf{x}_{n}\in\mathbb{R}^{m}$ and the SSM is applied elementwise on each feature. $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\in\mathbb{R}^{k}$, the recurrent form of a linear SSM layer defines a new sequence of states $\mathbf{h}_{1},\ldots,\mathbf{h}_{n}\in\mathbb{R}^{d}$ using projections $\bar{\mathbf{A}}_{i}\in\mathbb{R}^{d\times d}$ and $\bar{\mathbf{B}}_{i}\in\mathbb{R}^{d\times k}$, which can themselves depend on $\mathbf{x}_{i}$. For each $1\leq i\leq n$,
$\mathbf{h}_{i}=\bar{\mathbf{A}}_{i}\mathbf{h}_{i-1}+\bar{\mathbf{B}}_{i}\mathbf{x}_{i}.$ (Recurrent form)
The convolutional form of the SSM layer defines the same [^3] The two forms express the same function over $\mathbb{R}$ or any other distributive datatype. Over floating points (Section 2.2), they are not guaranteed to be the same, but we must assume the error is negligible for them to be well-defined and usable in practice. $\mathbf{h}_{1},\ldots,\mathbf{h}_{n}$ computed differently as a summation:
$\mathbf{h}_{i}=\sum_{j=1}^{i}\left(\prod_{k=j+1}^{i}\bar{\mathbf{A}}_{k}\right)\bar{\mathbf{B}}_{j}\mathbf{x}_{j}.$ (Convolutional form)
The layer outputs $\mathbf{y}_{i}=\mathbf{C}_{i}\mathbf{h}_{i}+\mathbf{D}_{i}\mathbf{x}_{i}\in\mathbb{R}^{k}$, where $\mathbf{C}_{i}\in\mathbb{R}^{k\times d}$ and $\mathbf{D}_{i}\in\mathbb{R}^{k\times k}$ depend on $\mathbf{x}_{i}$.
Two common cases of this layer are when $\bar{\mathbf{A}}_{i}$ does not depend on the input (“non-gated”; Section 4.2) and when $\bar{\mathbf{A}}_{i}$ is diagonal (Section 4.3). In both of these cases, we will show that the SSM can be simulated in $\mathsf{TC}^{0}$.
A generalized linear SSM is made up of multiple such layers, with a linear projection and a non-linearity applied after every layer (Rush & Karamcheti, 2022). Layer-norm can also be applied, either before or after the layer.
Practical Details.
In S4 and related SSMs, Definition 2.1 is applied elementwise ($k=1$) across all $m$ elements of the previous layer output (Gu et al., 2022a). In practice, the weight matrix initialization is crucial for training. Our expressivity results (Theorems 4.2 and 4.4) apply for any generalized linear SSM (including S4 and S6), independent of initialization. In contrast to S4 and S6, H3 (Fu et al., 2023) does not meet Definition 2.1 because the context is not represented by a single vector. Rather, it resembles a transformer with SSM components.
@@ 2.2 Numeric Datatype
Our main results (Theorems 4.2 and 4.4) will go through for any datatype $\mathbb{D}$ for which the following operations are efficiently parallel-computable (i.e., are in the complexity class $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, to be defined shortly in Section 2.3):
@@
    Matrix powering, i.e., computing the $n$-th power of a fixed-size $d\times d$ matrix over $\mathbb{D}$
When $\mathbb{D}$ is any finite-precision datatype, i.e., has a fixed number of bits available (e.g., 16 or 64), then these operations are easily seen to be in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. As Merrill & Sabharwal (2023b) argue, however, finite-precision datatypes severely limit the expressivity of neural architectures from a formal perspective (e.g., finite-precision transformers cannot represent uniform attention), motivating the use of parameterized datatypes that can (approximately) represent any number with a sufficiently large parameter. Interestingly, when $\mathbb{D}$ is the datatype of $n$-bit integers, all of the above operations are known to be in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (Hesse, 2001; Mereghetti & Palano, 2000). Realistically, however, neural model implementations use floating point numbers with much fewer than $n$ bits. Following Merrill & Sabharwal (2023b), we use the log-precision floating point model, i.e., $c\log n$ bit floats where $c$ is some fixed constant (see Appendix A for a formal definition). Merrill & Sabharwal (2023a) showed that iterated addition over log-precision floats is in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$. We extend the arguments of Hesse (2001) and Mereghetti & Palano (2000) to show that iterated product and matrix powering over log-precision floats are also in $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ (see Appendix A).
@@ 2.3 Limits of Transformers via Circuit Complexity
A line of recent work has used circuit complexity and logic formalisms to identify expressivity limitations of transformers on reasoning problems (Angluin et al., 2023; Merrill & Sabharwal, 2023a; Liu et al., 2023; Chiang et al., 2023; Merrill & Sabharwal, 2023b; Hao et al., 2022); see Strobl et al., 2024 for a survey. In particular, Merrill & Sabharwal (2023a) showed transformers can only solve problems in the complexity class $\mathsf{TC}^{0}$, which is the set of problems that can be recognized by constant-depth, polynomial-size threshold circuit families. Such circuits, in addition to having standard AND, OR, and NOT gates (of arbitrary fan-in), can also use threshold gates that output $1$ iff at least $k$ of the inputs are $1$, where $k$ is a parameter of the gate. Informally, $\mathsf{TC}^{0}$ can be thought of as the class of problems that can be solved with extremely parallel (constant-depth) computation. [^4] We use $\mathsf{TC}^{0}$ to mean $\mathsf{L}$-uniform $\mathsf{TC}^{0}$, meaning the circuit family is constructible by a Turing machine that runs in space logarithmic in the size of the input (cf. Merrill & Sabharwal, 2023a; Strobl et al., 2024). We believe our results could be extended from $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ to $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ using techniques similar to Merrill & Sabharwal (2023b) for composing $\mathsf{TC}^{0}$ circuits in a way that preserves $\mathsf{DLOGTIME}$ uniformity.